### v1

# A Unified Framework for Direction and Diffuseness Estimation Using Tight-Frame Microphone Arrays

Abstract
This work presents a unified framework for estimating both sound-field direction and diffuseness using practical microphone arrays with different spatial configurations. Building on covariance-based diffuseness models, we formulate a velocity-only covariance approach that enables consistent diffuseness evaluation across heterogeneous array geometries without requiring mode whitening or spherical-harmonic decomposition. Three array types — an A-format array, a rigid-sphere array, and a newly proposed tight-frame array — are modeled and compared through both simulations and measurement-based experiments. The results show that the tight-frame configuration achieves near-isotropic directional sampling and reproduces diffuseness characteristics comparable to those of higher-order spherical arrays, while maintaining a compact physical structure. We further examine the accuracy of direction-of-arrival estimation based on acoustic intensity within the same framework. These findings connect theoretical diffuseness analysis with implementable array designs and support the development of robust, broadband methods for spatial-sound-field characterization.
Submitted to the Journal of the Acoustical Society of America (under review).
I Introduction
A diffuse sound field is an ideal state in which both sound pressure and particle velocity are spatially and directionally isotropic. This assumption underpins the classical reverberation theories of Sabine and Eyring [1, 2, 3] and remains central in architectural acoustics [4, 5]. Quantitative evaluation of “diffuseness” thus plays a key role—not only for objectively assessing sound fields but also in applications such as room design, spatial reproduction, and auditory analysis.
Historically, early studies derived diffuseness indicators from spatial fluctuations of reverberation time [6] and from statistical limits in frequency-domain descriptions of sound fields [7]. Theoretical advances introduced diffusion factors to describe energy homogeneity [8] and correlation-based coefficients [9]. The spatial correlation of sound pressure in a diffuse field was shown to follow a sinc-type function [10].
In parallel, Thiele visualized reflection directions and strengths in concert halls, linking diffuseness with perceptual attributes such as clarity [11, 12, 13]. Following this, Gover and colleagues [14, 15] developed spherical microphone-array systems capable of analyzing the directional and spatial variations of reverberant fields. By steering beams over multiple directions, they computed the variance of directional power to define “directional diffusion” and the anisotropy index, showing how diffuseness evolves over time as the reverberant field becomes more isotropic and later anisotropic. These studies established the idea that a diffuse field can be characterized by the uniformity of directional energy distribution, forming an experimental foundation for later diffuseness metrics.
From the late 1980s to 1990s, the energy-transport equation provided a physically grounded framework for describing energy propagation in rooms. Derived from Boltzmann’s transport theory, it expresses temporal variation of acoustic energy density as a balance between flux and loss terms. Applying a Fourier-type approximation to the flux term yields the acoustic diffusion equation, which models energy distribution continuously in enclosed spaces. Picaut et al. [16] developed a numerical model based on this equation, later extended by Valeau et al. [17] to include anisotropy and boundary conditions. Hanyu [18] further incorporated wall scattering and absorption coefficients to characterize room diffuseness quantitatively. These developments form the theoretical core of modern room-diffusion analysis. Cox and D’Antonio [19] organized diffuser design and diffuse-field modeling into practical methodologies, and the ISO 17497 series [20, 21] standardized measurement and evaluation methods internationally.
In the 2000s, diffuseness quantification in time and space advanced rapidly. The First Order Ambisonics (FOA) based intensity-energy ratio (I/E) method by Merimaa and Pulkki [22, 23] characterized and reproduced sound fields by their diffuseness. Acoustic intensity was obtained from impulse responses, allowing estimation of diffuseness and major reflection directions, later reproduced using Vector Base Amplitude Panning (VBAP) [24] in Spatial Impulse Response Rendering (SIRR). The method was extended to continuous signals as Directional Audio Coding (DirAC) [25, 26], maintaining the same intensity-diffuseness principle. Since then, the method has evolved to estimate prominent reflections from the generalized cross-correlation of signals acquired with multiple microphones, named Spatial Decomposition Method, (SDM) [27], and more recently, applications in the ultrasonic domain [28] have also been reported.
Diffuseness estimation then became closely linked with spherical microphone arrays and spherical-harmonic analysis [29]. Epain and Jin [30] made a significant contribution with the covariance matrix eigenvalue diffuseness estimation, abbreviated as COMEDIE, method, which interprets diffuseness as a function of the eigenvalue distribution of spherical-harmonic signal covariance. In addition, they formulated a set of benchmark scenarios—such as single plane-wave, mixed diffuse-plus-direct, and interference fields—that have since served as reference frameworks for validating diffuseness estimators.
Following the covariance-based formulation by Epain and Jin[30], Politis and Pulkki[32, 31] extended this concept to spatially weighted energy analysis and parametric spatial audio frameworks. Their work unified the estimation of intensity, energy density, and diffuseness in the HOA domain and demonstrated its applicability to spaced microphone arrays and multichannel rendering. More recently, McCormack et al. [33] investigated the perceptual effects of spherical order, dedicated diffuse rendering, and frequency resolution in higher-order Spatial Impulse Response Rendering (SIRR), highlighting that perceived spaciousness strongly depends on the accuracy of diffuse-field modeling.
Throughout this history, the link between diffuseness estimation and intensity measurement has been widely recognized, as clearly demonstrated by the I/E method. The p-p probe method summarized by Fahy [34, 35] estimates local energy flow and reflection distributions. Guy and Abdou [36] visualized reflections and energy flow, while de Bree et al. [37, 38] introduced pressulre-velocity sensors. Zhang et al. [39] applied this p-u probes to near-field acoustic holography. Hoshi and Hanyu [40, 41] proposed estimating intensity using paired cardioid microphones. These studies advanced an integrated view of intensity and diffuseness, linking energetic analysis with spatial measurement.
In the present study, we aim to bridge theoretical diffuseness analysis and practical microphone measurements. A multichannel microphone array is employed to estimate broadband acoustic intensity simultaneously. By adopting uniform directional sampling based on a tight-frame array, the proposed method achieves both a higher usable frequency and improved stability. Combined with impulse-response measurements, it enables high-resolution evaluation of spatial energy distributions, including reflections.
While the proposed framework enables stable and broadband evaluation, it also reveals a fundamental limitation of conventional approaches: Conventional methods often reduce diffuseness to a single scalar, neglecting directional variation. In real environments, however, diffuseness depends on surface reflection and absorption distributions, making directional diffuseness essential. Its prototype appears in Meyer’s “hedgehog” visualization of directional energy using a parabolic microphone [12, 13]. This approach evolved into the SDM [27], psychoacoustic studies by Amengual et al. [42], and visualization work by Pätynen et al. [43], forming the basis for direction-dependent diffuseness analysis.
Building upon these theoretical foundations, the present study extends diffuseness estimation to practical microphone arrays with several different perspectives:
Microphone array based on tight-frame theory and spatial sampling — A tight frame is an overcomplete basis that enables redundant yet stable reconstruction [44]. On the sphere, a spherical $(t,t)$-design with $t=1$ is equivalent to a tight frame [45], providing a foundation for isotropic sampling in spherical processing. Recent work extends this concept to spherical framelets [46], optimal finite-dimensional frames[47]. Building on these, the present study implements uniform directional sampling based on a tight-frame array for diffuseness and intensity measurement.
Velocity-only covariance framework — a formulation that allows consistent diffuseness evaluation without relying on mode decomposition or ideal spherical harmonics, thus enabling unified analysis across heterogeneous array geometries.
Systematic comparison across indices and array configurations — Three array types: A-format, Tight Frame, and Rigid Sphere — are compared under identical conditions to evaluate consistency, frequency dependence, and the effects of spherical-harmonic order and whitening. Emphasis is placed on the trade-off between the numerical instability of high-order spherical harmonic (HOA) analysis [30] and the implementation stability of first-order (FOA) approaches. Overall, these analyses aim to bridge theoretical validity and practical reproducibility, establishing a unified framework for diffuseness evaluation.
Figure 1: Coordinate system used in this study.
II Microphone arrays for diffuseness assessment
Diffuseness estimation mainly relies on acoustic intensity and energy density, as well as on eigenvalue analyses of covariance matrices of spherical-harmonic (SH) components. Since both contain spatial and directional information, accuracy and effective bandwidth depend strongly on the geometry of the microphone array. Commonly used devices include p-p probes, p-u sensors, paired cardioid microphones, and spherical arrays such as the Eigenmikea). In particular, A-format microphones and rigid-sphere arrays are widely used for practical analysis.
An A-format microphone often conceals the directional responses of its capsule, hindering quantitative validation. Conversely, rigid-sphere arrays enable high-order SH expansion but require many channels, complex calibration, and suffer from numerical instability at high orders. Hence, there is a growing need for arrays that achieve uniform spatial sampling while remaining simple and robust.
In this study, we model both the A-format and rigid-sphere microphone arrays and further propose a stably configured array referred to as a tight-frame arrangement, examining the characteristics of each. The coordinate system used in the following discussion is shown in Fig. 1.
Figure 2: Microphone arrangements used in this study. (a) Microphone arrangement modeled after the AMBEO VR Mic (A-format array). (b) Array model based on the Eigenmike, in which 64 omnidirectional microphones are uniformly distributed on a rigid-sphere baffle according to a Fibonacci lattice. (c) Proposed tight-frame microphone configuration, in which microphones are spaced every 45° in azimuth and ±45° in elevation.
II.1 A-format microphone (Afmt)
An A-format microphone uses four capsules in a regular tetrahedral configuration, from which the FOA components are derived for straightforward spatial analysis. This configuration, referred to as Afmt, inherently offers limited spatial resolution and reduced directional discrimination at high frequencies.
In the simulations, the A-format microphone is modeled after the AMBEO VR Mic b), consisting of four cardioid capsules oriented along the axes shown in Fig. 2(a) within the coordinate system of Fig. 1:
$\bm{d}_{1}$ $=\frac{1}{\sqrt{3}}(+1,+1,+1)$ $:\;\;\text{FLU (Front Left Up)},$
$\bm{d}_{2}$ $=\frac{1}{\sqrt{3}}(+1,-1,-1)$ $:\;\;\text{FRD (Front Right Down)},$
$\bm{d}_{3}$ $=\frac{1}{\sqrt{3}}(-1,+1,-1)$ $:\;\;\text{BLD (Back Left Down)},$
$\bm{d}_{4}$ $=\frac{1}{\sqrt{3}}(-1,-1,+1)$ $:\;\;\text{BRU (Back Right Up)}.$
Let $D_{1}$ to $D_{4}$ be the complex amplitude of measured signals. They are converted into the B-format signals $(W,X,Y,Z)$ as
$\begin{bmatrix}W\\ X\\ Y\\ Z\end{bmatrix}=\frac{1}{2}\begin{bmatrix}+1&+1&+1&+1\\ +1&+1&-1&-1\\ +1&-1&+1&-1\\ +1&-1&-1&+1\end{bmatrix}\begin{bmatrix}D_{1}\\ D_{2}\\ D_{3}\\ D_{4}\end{bmatrix}$. (1)
The corresponding sound pressure and particle velocity are
$p=W,\quad\bm{u}=-\frac{\sqrt{3}}{Z_{0}}\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}$ (2)
where $Z_{0}$ is the characteristic acoustic impedance of the medium. From these quantities, both intensity and the total energy can be derived as
$\bm{I}=\frac{1}{2}\Re\{\,p\,\bm{u}^{\ast}\,\},\quad E=\frac{|p|^{2}}{4\rho_{0}c^{2}}+\frac{\rho_{0}}{4}|\bm{u}|^{2},$ (3)
where $\Re\{\cdot\}$ denotes the real part and superscript $(\cdot)^{*}$ denotes complex conjugation. Also, $\rho_{0}$ and $c$ represent the mass density of the medium and the sound speed, respectively. The ratio $|\bm{I}|/(cE)$ serves as the diffuseness index in the I/E method.
II.2 Rigid-sphere baffle microphone array (Fibo64)
Spherical arrays such as the Eigenmike place microphones on a rigid sphere and use SH analysis to decompose the sound field into higher-order components, enabling analysis of both overall diffuseness and local anisotropy. Practical challenges remain:
  • Accurate modeling of the sphere and baffle acoustics is required.
  • Numerical stability degrades at high orders.
  • Dense sampling increases calibration and cost.
Thus, despite their theoretical power, spherical arrays face trade-offs between spatial uniformity and implementation simplicity.
In this study, a 64-channel array of omnidirectional microphones uniformly distributed on a rigid sphere (radius 0.042 m) is simulated. From the simulated pressures, sound pressure, particle velocity, and HOA coefficients are computed following Rafaely[29].
Sound pressure on a rigid sphere
The pressure on a rigid sphere of radius $r$ can be expanded as
$p(\Omega,k)=\sum_{n,m}a_{nm}(k)\,b_{n}(kr)\,Y_{nm}(\Omega),$ (4)
where $b_{n}(kr)$ is the radial filter corresponding to the rigid-boundary condition for the wavenumber $k$. Each microphone measures the pressure at direction $\Omega_{q}$.
Let $\bm{p}=[p_{1},p_{2},\dots,p_{Q}]^{\!\top}$ denote the pressures measured at $Q$ positions $\{\Omega_{q}\}$. The discrete model is then expressed as
$\bm{p}=\bm{Y}\,\bm{B}\,\bm{a},$ (5)
where $\bm{Y}$ is the spherical harmonic (SH) matrix, $\bm{B}$ is the diagonal matrix of radial filters, and $\bm{a}$ is the vector of higher-order coefficients.
Estimation of HOA coefficients
The estimation of HOA coefficients $\widehat{\bm{a}}$ are then estimated in the least-squares sense as
$\widehat{\bm{a}}=\bm{B}^{-1}\bm{Y}^{\dagger}\bm{p},$ (6)
where $\bm{Y}^{\dagger}$ is the pseudoinverse of $\bm{Y}$. In the practical calculation, the Tikhonov regularization is introduced with the small regularization parameter, say, $10^{-4}$.
The obtained coefficients with $n=0$ and $n=1$ correspond to pressure and velocity:
$p=\frac{a_{00}}{\sqrt{4\pi}},\quad\bm{u}=\frac{1}{3i\rho_{0}c}\sqrt{\frac{3}{4\pi}}\begin{bmatrix}a_{1x}\\ a_{1y}\\ a_{1z}\end{bmatrix}$. (7)
From the pressure and the velocities above, the intensity and the energy can be calculated as Eq. (3). These four coefficients $\{a_{00},a_{1x},a_{1y},a_{1z}\}$ constitute the FOA components. The higher-order coefficients are used as necessary.
For this array (64 channels, radius $r$ = 0.042 m), the spherical harmonic expansion was limited to the fixed order $N_{\mathrm{max}}=4$. This corresponds approximately to $kr\approx 4$, i.e., a validity limit near $f=\frac{4c}{2\pi r}\approx$5.2\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$\$ for $r=$42\text{\,}\mathrm{m}\mathrm{m}$\$. Higher frequencies were also analyzed for completeness, but the HOA representation above this limit may include spatial aliasing effects. The fixed-order setting was chosen to ensure a consistent comparison across arrays and to avoid overfitting the mode expansion to higher frequencies. This 64-point Fibonacci-based equal-area distribution, modeled as a proxy for the Eigenmike, was referred to as Fibo64, hereafter.
II.3 Tight frame microphone array (TF24)
We propose a new array using multiple highly directional microphones arranged isotropically on a sphere based on a tight-frame distribution, referred to as TF24. A finite set of such points uniformly approximates the sphere while maintaining reconstruction stability. This microphone configuration was initially devised for measuring directional impulse responses intended for sound field reproduction[48]. Preliminary trials were conducted to estimate acoustic intensity using this setup, which subsequently led to the more detailed investigation presented in this study.
Advantages of the proposed array include:
  • Reduced angular-dependent errors through uniform sampling
  • Wideband operation with fewer channels
  • Easier calibration using well-characterized microphones
  • Possibility of sequential measurements by rotating a single microphone, in the case of impulse response measurement
Let the measured directional quantities, say, intensities, form $\bm{I}_{\text{meas}(N)}$, and let $\bm{R}$ be the $N\times 3$ matrix of unit direction vectors $\bm{r}_{i}$:
$\bm{I}_{\text{meas}(N)}=\begin{bmatrix}I_{r_{1}}\\ I_{r_{2}}\\ \vdots\\ I_{r_{N}}\end{bmatrix},\quad\bm{R}=\begin{bmatrix}r_{1,x}&r_{1,y}&r_{1,z}\\ r_{2,x}&r_{2,y}&r_{2,z}\\ \vdots&\vdots&\vdots\\ r_{N,x}&r_{N,y}&r_{N,z}\end{bmatrix}$. (8)
The Cartesian intensity estimate is
$\bm{I}_{\text{meas}(xyz)}=\bm{R}^{\!\top}\bm{I}_{\text{meas}(N)}.$ (9)
For the true intensity $\bm{I}_{(xyz)}$, its projection onto $N$ measured directions is
$\bm{I}_{(N)}=\bm{R}\bm{I}_{(xyz)},$ (10)
and reconstruction via the pseudoinverse yields
$\bm{I}_{(xyz)}=\bm{R}^{\dagger}\bm{I}_{(N)}=(\bm{R}^{\!\top}\bm{R})^{-1}\bm{R}^{\!\top}\bm{I}_{(N)}.$ (11)
For a tight frame, $\bm{R}^{\!\top}\bm{R}=A\bm{I}$ is assured and this relationship gives
$\bm{I}_{(xyz)}=\frac{1}{A}\bm{R}^{\!\top}\bm{I}_{(N)}.$ (12)
Thus, the projection sum in Eq. (9) yields a scaled Cartesian intensity vector.
Microphones and configuration.
Twelve directional axes are defined using twenty-four microphones at three zenith angles ($\theta=45^{\circ},\;90^{\circ},\;135^{\circ}$) and eight azimuths separated by $45^{\circ}$. Each opposing pair forms one velocity axis. The direction matrix $\bm{R}$ is
$\bm{R}=\begin{bmatrix}0&1/\sqrt{2}&1/\sqrt{2}\\ 1/2&1/2&1/\sqrt{2}\\ 1/\sqrt{2}&0&1/\sqrt{2}\\ 1/2&-1/2&1/\sqrt{2}\\ 0&-1/\sqrt{2}&1/\sqrt{2}\\ -1/2&-1/2&1/\sqrt{2}\\ -1/\sqrt{2}&0&1/\sqrt{2}\\ -1/2&1/2&1/\sqrt{2}\\ 0&1&0\\ 1/\sqrt{2}&1/\sqrt{2}&0\\ 1&0&0\\ 1/\sqrt{2}&-1/\sqrt{2}&0\end{bmatrix}$, (13)
and $\bm{R}^{\!\top}\bm{R}=4\bm{I}$ in this case.
Directional intensity, energy, and diffuseness
Microphone pairs aligned along twelve directions compute pseudo-intensity via the cardioid-pair (C-C) principle [40, 41], here adapted for narrow-directional microphones as the pseudo C-C method.
As illustrated in Fig. 3, let $M^{+}_{\bm{r}_{i}}$ and $M^{-}_{\bm{r}_{i}}$ denote the complex amplitudes of the signals at a given frequency measured by the microphones located at the positive and negative ends of the $\bm{r}_{i}$ axis, respectively, for $i=1,\ldots,12$.
The directional response of each microphone, $d(\theta)$, where $\theta$ denotes the angle from its main axis, is approximated by a polynomial of $\cos\theta$ as
$d(\theta)=\sum_{n=0}^{N}a_{n}\cos^{n}\theta.$ (14)
The case of $N=1$ with $a_{0}=0.5$ and $a_{1}=0.5$ corresponds to an ideal cardioid pattern.
When two microphones are aligned along the same axis but oriented in opposite directions, their directional responses can be expressed as $d(\theta)$ and $d(\pi-\theta)$, respectively. The sum and difference of their complex amplitudes,
$M^{+}_{\bm{r}_{i}}\pm M^{-}_{\bm{r}_{i}},$
then form symmetric and antisymmetric combinations of these directional responses. The sum component primarily exhibits an omnidirectional characteristic, approximating the local sound pressure, whereas the difference component resembles a figure-of-eight pattern, corresponding to the particle velocity component along $\bm{r}_{i}$. Hence, the pairwise sum and difference of oppositely oriented directional microphones provide a practical means of approximating the local pressure and the corresponding particle-velocity component.
The resulting quantities are defined as
$p_{\bm{r}_{i}}$ $=M^{+}_{\bm{r}_{i}}+M^{-}_{\bm{r}_{i}},$ (15)
$u_{\bm{r}_{i}}$ $=\frac{M^{+}_{\bm{r}_{i}}-M^{-}_{\bm{r}_{i}}}{Z_{0}},$ (16)
$I_{\bm{r}_{i}}$ $=\frac{|M^{+}_{\bm{r}_{i}}|^{2}-|M^{-}_{\bm{r}_{i}}|^{2}}{2Z_{0}},$ (17)
$E_{\bm{r}_{i}}$ $=\frac{|p_{\bm{r}_{i}}|^{2}}{4\rho_{0}c^{2}}+\frac{\rho_{0}}{4}|u_{\bm{r}_{i}}|^{2}.$ (18)
Figure 3: Schematic diagram of the signals used to construct the pseudo C–C method in the TF24 system. Twelve directional pairs are formed from 24 microphones.
In practice, the directional elements used in the array do not have strictly cardioid directivity but rather exhibit sharper lobes depending on frequency. Therefore, the pressure obtained from their summation and the velocity inferred from their differential output cannot be regarded as the exact physical quantities. Instead, they represent operational approximations of the local sound pressure and particle velocity, derived under the assumption of complementary directivity pairs. The corresponding acoustic intensity and energy density are thus defined in this approximate sense and referred to here as pseudo-intensity and pseudo-energy density. Representative examples of these directivity patterns are presented in the next section.
Similarly, the directional pseudo-diffuseness index based on the I/E concept is defined as
$\Psi_{\bm{r}_{i}}=1-\frac{|I_{\bm{r}_{i}}|}{E_{\bm{r}_{i}}c},$ (19)
and its weighted extension is discussed in the following section.
III Benchmark problems for numerical simulation
Unlike previous studies that evaluated diffuseness primarily within the spherical harmonic domain under ideal conditions, the present work reformulates the covariance analysis in the first-order velocity space and verifies its performance with physically realizable microphone geometries. By integrating Epain and Jin’s benchmark paradigm[30] with a tight-frame sampling approach, this study provides a unified yet experimentally grounded framework for diffuseness characterization.
III.1 Overview of simulation framework
This section defines benchmark problems to compare the diffuseness-evaluation methods for the three array configurations introduced in the previous section (Afmt, Fibo64, and TF24). Each benchmark specifies a virtual acoustic scene with known structure; from the simulated microphone pressures, we compute diffuseness and assess accuracy, stability, and frequency dependence.
The Afmt array was modeled after the AMBEO VR Mic, and the capsule directivities were approximated by a cardioid model consistent with DPA 2012. For the TF24 array, we used a highly directional model consistent with DPA 4017. In both cases, directivity patterns were measured in an anechoic chamber at $3^{\circ}$ or $5^{\circ}$ resolution and fitted as a polynomial in $\cos\theta$ (power-series form) up to 8th order; these fits were used in simulation (Fig. 4).
Figure 4: Measured (solid lines) and modeled (broken lines) directivities of the (a) DPA2012 cardioid microphone assumed in Afmt array and (b) DPA 4017 narrow-directional microphone used in the TF24 array. To avoid visual clutter, the polar plots show the responses for 250 Hz to 1 kHz on the left half and for 2 kHz to 8 kHz on the right half. In the calculations presented in the text, however, a wider frequency range from 63 Hz to 16 kHz was analyzed in octave bands.
Figure 5: Example of modeled directional responses derived from the microphone directivity polynomial of Eq. (21) at 1 kHz. The outer line shows the pseudo-omnidirectional response (sum signal), and the inner line shows the pseudo-figure-of-eight response (difference signal). Both were obtained from paired microphones placed in opposite directions in the TF24 array.
Array geometries were set as follows (see Fig. 2): the A-format capsules are located 6 mm from the origin [Fig. 2(a)]; the rigid-sphere array uses a radius of 42 mm [Fig. 2(b)]. The spacing between microphones in the TF24 array was set to 20 mm; that is, the distance from the origin in [Fig. 2(c)] is 10 mm. Although this dimension may seem unrealistic for highly directional microphones approximately 200 mm in length, it represents a practical configuration assuming sequential impulse-response measurements by mechanically rotating a single microphone.
Simulations are performed in the frequency domain. Octave bands from $63\text{\,}\mathrm{H}\mathrm{z}$ to $16\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$ were analyzed. Sources are either a single plane wave or a sum of multiple plane waves. To emulate band-limited noise, we draw 100 frequencies uniformly in log scale within a one-octave band and sum sinusoids with random amplitude and phase.
Pressure at each microphone
The complex pressure at the $m$-th microphone for frequency $f$ (wavenumber $k=2\pi f/c$) due to $N_{s}$ plane waves incident from directions $\hat{\nu}_{i}$ is expressed as
$p_{m}(f)=S_{m}(f)\sum_{i=1}^{N_{s}}D_{m}\!\bigl(\hat{\nu}_{i},f;\hat{d}_{m}\bigr)A_{i}\,e^{\,jk\,\bm{r}_{m}\cdot\hat{\nu}_{i}+j\varphi_{i}},$ (20)
where $\bm{r}_{m}$ is the position vector of the $m$-th microphone, $\hat{\nu}_{i}$ is the unit vector of the $i$-th plane-wave propagation direction, $A_{i}$ and $\varphi_{i}$ are its amplitude and phase, respectively, $S_{m}(f)$ denotes the on-axis sensitivity of the microphone, and $D_{m}(\hat{\nu}_{i},f;\hat{d}_{m})$ represents the directional response of the microphone whose principal axis points along the unit vector $\hat{d}_{m}$. The exponential term $e^{jk\bm{r}_{m}\!\cdot\hat{\nu}_{i}}$ accounts for the phase delay due to the microphone position.
For the present simulations, each directivity pattern was modeled as a polynomial function of $\cos\theta$, where $\theta$ is the angle between $\hat{d}_{m}$ and $\hat{\nu}$:
[…]
This fitted model corresponds to the broken lines shown in Fig. 4.
As an example, Fig. 5 illustrates the modeled directivities obtained from the sum and difference of opposite microphone pairs based on the polynomial model in Eq. (21) at 1 kHz. The sum response approximates a pseudo-omnidirectional pattern, while the difference response approximates a pseudo-figure-of-eight pattern. For the Fibo64 rigid-sphere model with omnidirectional elements, $D_{m}=1$; directivity/baffle effects are captured by the radial filters $b_{n}(ka)$ in the Spherical Harmonic expansion. Thus, after computing $p_{m}(f)$ via (20), HOA coefficients are estimated from the SH model as described in Sec. II.
Figure 6: Conceptual diagrams of the incident directions of plane waves used in the benchmark simulations: (a) single plane-wave incidence from various directions; (b) diffuse components arriving randomly over the entire solid angle of $4\pi$, and beam components arriving within a confined solid angle corresponding to 0.5 % of the total sphere. The center of the beam was set at 3° in azimuth and 87° in zenith angles. (c) An interference field formed by two plane waves. One (the primary wave) was incident from the zenith direction ($\theta$ = 0°), while the other secondary wave was incident from variable zenith angles ranging from 0° to 180°, producing an interference pattern used to evaluate the diffuseness under such conditions.
III.2 Benchmark case definitions
As shown in Fig. 6, three representative cases are assumed for numerical simulation.
Case 1: Single plane-wave incidence.
A single plane wave arrives from a given direction; the theoretical diffuseness is $\Psi=0$. We sweep both elevation and azimuth over the sphere to test directional robustness and verify how accurately intensity-based direction-of-arrival can be recovered. For Case 1, 72 azimuthal directions (0 to 355) and 35 zenith angles (5 to 175), both at 5 intervals, were simulated, resulting in 2520 incidence directions in total. For each case, the mean diffuseness and the error in arrival-angle detection based on the acoustic intensity were evaluated to assess the accuracy of sound-arrival direction estimation for each array.
Case 2: Plane-wave beam + diffuse mixture.
As illustrated in Fig. 6, we mix a beam arriving from a narrow solid angle (here $0.5\%$ of $4\pi$ sr) with diffuse components uniformly distributed over $4\pi$ sr. We define the beam-to-total energy ratio
$\eta=\frac{E_{\mathrm{beam}}}{E_{\mathrm{beam}}+E_{\mathrm{diffuse}}},$ (22)
where $E_{\mathrm{beam}}$ is the energy of beam, and $E_{\mathrm{diffuse}}$ is the energy of diffuse components. The $\eta$ was varied from 0 to 1. Both the diffuse and beam components use 100,000 rays with random amplitude and phase; for band-limited noise, we draw 100 log-spaced frequencies within a 1-octave band at each trial. We examine whether the estimated $\Psi$ is stable and how linearly it tracks $\eta$.
Case 3: Reactive field (interference).
In this case, an interference field was simulated by superposing two plane waves: a primary wave incident from the zenith direction and a secondary wave arriving from variable zenith angles between 0° and 180°. For each configuration, 1,000 realizations were generated by randomly varying the phase (within $\pm 2\pi$) and amplitude (within $\pm 3$ dB) of the secondary wave. The acoustic intensity, energy density, and covariance matrix were computed for each realization, and the final values were obtained by averaging these quantities over all realizations to represent the mean behavior of the interference field.
III.3 Evaluation metrics
Given the array pressures, the following diffuseness indices are calculated.
Intensity–Energy ratio $\Psi_{\mathrm{IE}},\>\Psi_{\mathrm{AVE}}$
From pressure $p$ and velocity $\bm{u}$ (via A-format or HOA/FOA processing) in the cases of Afmt and Fibo64, the diffuseness $\Psi_{\mathrm{IE}}$ is defined as
$\Psi_{\mathrm{IE}}(f)=1-\frac{|\bm{I}(f)|}{c\,E(f)}.$ (23)
$\Psi_{\mathrm{IE}}\!\approx\!0$ indicates a traveling-wave field; $\Psi_{\mathrm{IE}}\!\approx\!1$ indicates a diffuse field.
For TF24 case, we compute directional $\Psi_{r_{i}}$ defined as Eq. (19). In directions where only uncorrelated background noise is observed, the estimated diffuseness tends to assume relatively high values. However, such regions do not correspond to acoustically significant signals. This observation suggests the necessity of introducing directional weighting.
One possible choice is to weight the diffuseness by the energy associated with each direction. After several preliminary trials, weighting by the squared magnitude of the particle velocity was adopted, as it exhibited higher sensitivity to spatial variations in the sound field. Fig. 7 illustrates an example of the directional energy distribution and the corresponding weighting factors based on the squared particle velocity, computed for Case 2. The incident beam is assumed to arrive predominantly from the 9th direction, and it can be observed that the velocity-based weighting responds subtly but consistently to this directional bias.
$\Psi_{\mathrm{AVE}}=\displaystyle\frac{\displaystyle\sum_{i=1}^{12}\Psi_{\bm{r}_{i}}\>|u_{\bm{r}_{i}}|^{2}}{\displaystyle\sum_{i=1}^{12}\>|u_{\bm{r}_{i}}|^{2}}.$ (24)
The appropriateness of this weighting scheme remains a subject for further investigation; nevertheless, several representative examples are presented below. The possible disadvantages associated with this choice are also discussed in the Discussion section.
Figure 7: Example of weighting applied to the directional diffuseness: results for Case 2 with a beam ratio of $\eta=0.95$ at 2 kHz.
Eigenvalue-based diffuseness ($\PsiPR, Ψ\PsiCOM$)
To capture spatial uniformity, covariance-based measures are employed. In conventional approaches, a 4×4 covariance matrix is constructed from the sound pressure and three velocity components, i.e., FOA components as
$\bm{C}_{\mathrm{pu}}=\langle[p,\,u_{x},\,u_{y},\,u_{z}]^{H}[p,\,u_{x},\,u_{y},\,u_{z}]\rangle,$
where $u_{x},\;u_{y},\;u_{z}$ are the $x,y,z$ directional components of velocity, and $\bm{C}_{\mathrm{pu}}$ represents the full energy coupling between pressure and particle velocity.
In the present study, however, we emphasize a velocity-only route and form
$\bm{C}_{\mathrm{uu}}=\langle\bm{u}(f)\,\bm{u}^{H}(f)\rangle,$ (25)
which directly captures the spatial anisotropy of the local particle-velocity distribution without explicit reference to the pressure term.
This choice is first motivated by the characteristics of the Tight frame array. Since each microphone in the TF24 configuration is a highly directional (DPA4017 in our case) rather than an ideal cardioid, the sum of each microphone pair does not accurately reproduce the omnidirectional pressure component.
Consequently, defining diffuseness in terms of pressure-velocity coherence becomes less reliable in TF24. By constructing the covariance matrix solely from the velocity vectors, we can evaluate the spatial uniformity in a physically consistent and array-independent manner, thus ensuring the robustness of diffuseness estimation even under non-ideal directivity conditions.
In the present study, the covariance matrix was constructed solely from the velocity components derived from the FOA signals, not only for theTF24 system but also for the Afmt and Fibo64 configurations. A further dominant motivation for this choice lies in the scaling issues that inevitably arise when handling both pressure and velocity simultaneously. When estimating diffuseness from eigenvalues, it is desirable to ensure not only amplitude scaling, such as SN3D normalization (e.g., Eq. (7)) but also detailed frequency response consistency. Achieving this would require measurements in a sound field that can be regarded as sufficiently diffuse and the computation of a corresponding whitening matrix. To avoid this complexity, the present work primarily adopts an approach based on velocity-only metrics.
In the TF24 system, the particle velocity components used to construct the covariance matrix $\bm{C}_{uu}$ are obtained by projecting the velocity vectors $u_{\bm{r}_{i}}$, measured in 12 directions, onto the global x,yx,\;y, and $z$ axes. This projection is performed by multiplying each directional component by the corresponding direction vector $\bm{R}$ as,
$\bm{u}=\begin{bmatrix}u_{x}\\ u_{y}\\ u_{z}\end{bmatrix}=\frac{1}{4}\bm{R}^{\!\top}\bm{u}_{\bm{r}},$ (26)
where $\bm{u}_{\bm{r}}=\left[u_{\bm{r}_{1}}u_{\bm{r}_{2}}\cdots u_{\bm{r}_{12}}\right]^{\!\top}$.
Let the eigenvalues of Eq. (25) be $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$. From these eigenvalues, several indices can be derived. In this paper, the following two diffuseness indices are calculated and examined:
[…]
    $\Psi_{\mathrm{PR}}=\frac{\displaystyle\left(\sum_{i=1}^{3}\lambda_{i}\right)^{2}}{\displaystyle 3\sum_{i=1}^{3}\lambda_{i}^{2}}.$ (27)
    This index is sensitive to how evenly the energy is distributed among the eigenvalues, taking a value close to $1/3$ for a single dominant mode and approaching 1 for a uniform distribution. Note that $\Psi_{\mathrm{PR}}\in[1/3,\,1]$ for three non-negative eigenvalues.
[…]
    $\Psi_{\mathrm{COM}}=1-\sqrt{\frac{3}{2}}\,\frac{\sigma_{\lambda}}{\overline{\lambda}},$ (28)
    where $\sigma_{\lambda}$ and $\overline{\lambda}$ denote the standard deviation and mean of $\{\lambda_{i}\}$, respectively. This provides a statistically stable and physically interpretable indicator of isotropy.
IV Results and discussion
Case 1: Diffuseness and sound-arrival direction estimation for a single plane wave
A single plane wave was simulated to impinge from all directions, and the I/E based diffuseness indices $\Psi_{\mathrm{IE}}$ for Afmt and Fibo64, and $\Psi_{\mathrm{AVE}}$ for TF24, and eigenvalue-based indices $\Psi_{\mathrm{PR}}$, and $\Psi_{\mathrm{COM}}$—were computed. In theory, all indices should yield $\Psi=0$, but small deviations were observed depending on the microphone array configuration, especially in the case of I/E based indices. Fig. 8 shows the results. The averaged value of diffuseness for all incident angles is shown at every frequency.
Figure 8: Case 1: Diffuseness indices based on the I/E method: for Afmt and Fibo64, the results show $\Psi_{\mathrm{IE}}$; for TF24, $\Psi_{\mathrm{AVE}}$ is presented.
The Afmt results exhibited relatively large values in the low-frequency region and above 4 kHz, while the TF24 array also showed slightly elevated values of approximately 0.2 at 16 kHz. In contrast, the Fibo64 array, which employs 64 microphones, yielded values close to zero across all frequency bands.
For the eigenvalue-based indices, although the absolute magnitudes differed across methods, all 2,520 incident directions yielded values of zero up to three decimal places when averaged over the nine octave bands from 63 Hz to 16 kHz.
Under the condition of single plane-wave incidence, the superiority of the eigenvalue-based analysis using the covariance matrix was demonstrated. Among the microphone arrays, the Fibo64 configuration exhibited the most stable results.
While the upper frequency limit of the Fibo-64 array can be roughly predicted from the spatial aliasing condition, the simulation results exhibit stable performance even above this range. This suggests that the quasi-uniform sampling of the Fibonacci lattice provides smoother angular coverage and mitigates high-order aliasing components that typically degrade array accuracy.
Directional accuracy of the active-intensity vector was also compared. The results are shown in Fig. 9. Although diffuseness and the angular error in intensity detection are entirely different quantities, the overall trend of the results is similar to that shown in Fig. 8. For the Afmt system, the estimation accuracy is low in both the low- and high-frequency ranges, and the frequency band in which the correct incident direction can be detected is limited to approximately 500 Hz to 2 kHz. In contrast, the Fibo64 system achieves almost error-free direction detection across the entire frequency range. The performance of TF24 lies between these two cases, showing an average detection error of about 4° at 8 kHz and slightly less than 6° at 16 kHz. Considering that these values represent the average error over 2520 incident directions, the overall performance can be regarded as practically sufficient.
Figure 9: Case 1: Mean detection error of the incident angle estimated from acoustic intensity obtained by each microphone array: Average detection error over 2520 incident directions.
Case 2: Linearity of diffuseness in plane-wave and diffuse mixture
Next, a mixture of a narrow beam and an isotropic diffuse field was simulated by varying the energy ratio $\eta$ between beam and diffuse components.
Figure 10 shows the results of the COMEDIE-based diffuseness index, $\Psi_{\mathrm{COM}}$, for the TF24 system. Although slight fluctuations can be observed in the high-frequency range, the results closely follow the theoretical value of $1-\eta$ from the low to mid frequencies. All indices exhibit an approximately linear relationship, and the correlation between $1-\eta$ and the diffuseness estimated in each frequency band is very high. Because the indices are therefore difficult to distinguish in direct comparison, we quantified their deviation more explicitly by calculating the absolute difference between the diffuseness value and $1-\eta$, and taking the maximum of this difference for each index. The results are shown in Fig. 11.
In Fig. 11, the results obtained using the intensity/energy (I/E)-based indices are shown as solid lines, while those based on the eigenvalue methods—particularly COMEDIE—are shown as dashed lines. The eigenvalue-based results show only small deviations across all frequency bands. A similar tendency to that seen in Figs. 8 and 9 is observed for the Afmt system, especially in the I/E-based indices, indicating that the frequency range with small estimation error is relatively narrow. In contrast, even for Afmt, the indices derived from the eigenvalues of the velocity covariance matrix exhibit comparatively small errors.
Figure 10: Case 2: Example of the variation in diffuseness when the beam ratio $\eta$ is varied from 0 to 1 in steps of 0.05. Results are obtained for the TF24 array using the eigenvalues of the covariance matrix and the COMEDIE, i.e., $\Psi_{\mathrm{COM}}$. The numbers in the figure indicate $\eta$; in the ideal case, the diffuseness corresponds to $1-\eta$.
Figure 11: Case 2: Maximum absolute error between the beam ratio $1-\eta$ and the diffuseness, calculated with I/E (solid lines) and eigenvalue examination; $\Psi_{\mathrm{COM}}$ (broken lines).
Case 3: Reactive (interference) field
Figure 12: Case 3: Diffuseness for the interference field formed by two plane waves at various incident angles at $1\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$. (a) Afmt, (b) Fibo64, and (c) TF24. The results for $\Psi_{\mathrm{IE}}$, $\Psi_{\mathrm{PR}}$, and $\Psi_{\mathrm{COM}}$ are shown. For reference, in (a), the results of $\Psi_{\mathrm{PR}}$ and $\Psi_{\mathrm{COM}}$ obtained using all four FOA components ($p$ and $\bm{u}$) are also presented. In (b), the results of $\Psi_{\mathrm{COM}}$ calculated from the HOA coefficients (up to 4th) are additionally shown.
Figure 12 shows the diffuseness results for Case 3, in which two plane waves interfere at various incidence angles. The results are for 1kHz. Panels (a), (b), and (c) correspond to the results for Afmt, Fibo64, and TF24, respectively. For all microphone arrays, both the intensity/energy-based index $\Psi_{\mathrm{IE}}$ and the eigenvalue-based indices $\Psi_{\mathrm{PR}}$ and $\Psi_{\mathrm{COM}}$ are presented. For reference, in (a) (Afmt), $\Psi_{\mathrm{PR/pu}}$ and $\Psi_{\mathrm{COM/pu}}$, which were computed from the eigenvalues of the $4\times 4$ covariance matrix including both pressure and velocity components after whitening, are also shown. In (b) (Fibo64), $\Psi_{\mathrm{COM/HOA}}$, obtained from the eigenvalues of the whitened covariance matrix of the HOA coefficients, is additionally presented.
A notable feature is that the $\Psi_{\mathrm{IE}}$ index approaches unity as the incidence angle between the two plane waves approaches 180°, that is, when two plane waves of the same frequency arrive from opposite directions and interfere. This high diffuseness value arises because the acoustic intensity vector becomes nearly zero due to destructive interference. However, from a physical standpoint, such a condition cannot be regarded as a truly diffuse sound field, since the waves propagate only along a single axis (the $z$-direction in this case).
In contrast, for the eigenvalue-based indices calculated using only the velocity components, the diffuseness reaches its maximum value of approximately 0.5 at an incidence angle of 90°, and then gradually decreases as the angle increases, converging to zero at 180°. This tendency is almost identical for both $\Psi_{\mathrm{PR}}$ and $\Psi_{\mathrm{COM}}$. (Although not shown here for simplicity, if other eigenvalue-based measures such as the maximum-eigenvalue ratio or entropy-based metrics are used, the resulting diffuseness values fluctuate around these trends.) In panel (a), when the pressure component is also included in the covariance matrix, the peak position shifts, and the diffuseness converges to a value around 0.2 to 0.3 at 180°.
For the Fibo64 array shown in panel (b), the diffuseness index based on the eigenvalues of the HOA coefficients, presented for reference, converges to an approximately constant value of 0.3 for incidence angles above about 45°. When the interference of two plane waves at 180° is interpreted as representing a single propagation direction, it is reasonable that the diffuseness converges to zero. This behavior can be interpreted as supporting the validity of estimating diffuseness from the eigenvalue analysis of the velocity covariance matrix alone.
Figure 13: Diffuseness calculated from waveforms obtained by mixing impulse responses measured in anechoic and reverberation rooms at various energy ratios. The energy ratio of the anechoic component was varied from 0 to 1 in 0.1 increments. The left panels show the results for the Afmt array, and the right panels show those for the TF24 array. The upper panels correspond to the intensity-energy (I/E) method, while the lower panels show the eigenvalue-based results using the COMEDIE index.
V Experimental verification
V.1 Concept
Finally, the trends obtained from the simulations were experimentally verified using measured data. Two sound fields were considered: an anechoic room and a reverberation room. As expected, diffuseness is nearly zero in the anechoic room and close to unity in the reverberation room. However, unlike the numerical simulations, neither environment can be regarded as perfectly ideal, and it is difficult to ensure that the measured results precisely match these theoretical values. Therefore, to enable a meaningful verification, we adopted an approach analogous to Case 2 of the benchmark simulations: the impulse responses measured in the anechoic and reverberation rooms were used to represent the “beam” and “diffuse” components, respectively. By mixing these two responses at arbitrary energy ratios, we examined the linearity between the prescribed ratio and the resulting diffuseness.
Specifically, for both sound fields, impulse responses were measured using the Afmt and TF24 microphone arrays. The energy within each frequency band was computed and combined according to the specified ratio, and the diffuseness was then calculated from the resulting signals. It should be noted that an Eigenmike was not available, and thus the case corresponding to the Fibo64 array was not experimentally validated.
V.2 Conditions
For the Afmt array, an AMBEO VR Mic was used, and for the TF24 array, DPA 4017 microphones were employed with a spacing of 20 mm between each microphone pair. As a preliminary measurement conducted in an anechoic room to determine the acoustic reference point of the shotgun microphone, impulse responses were measured twice at the same position: once with the microphone oriented toward the sound source, and once with it oriented perpendicularly. The time difference between the peaks of these two responses was used to estimate the reference point, which was found to be approximately 61 mm from the bottom of the microphone (amplifier section). Using this reference, the array was mounted on a rotatable jig (a camera tripod head) that allowed precise rotations in 45° increments around a point located 10 mm above the estimated center.
In the anechoic chamber, measurements were conducted at a distance of 2.5 m from the sound source. In the reverberation chamber (147 m3, RT $\sim$ 4.7 s), the distance to the source, which was directed toward a wall, was approximately 5 m.
V.3 Results
The results are shown in Fig. 13. The left panels correspond to the Afmt array, and the right panels to the TF24 array. The upper panels show the diffuseness obtained using the intensity-energy (I/E) method, while the lower panels present the diffuseness $\Psi_{\mathrm{COM}}$ calculated from the eigenvalues of the covariance matrix constructed only from the velocity components of the FOA signals. It is evident that the results for the Afmt array exhibit a larger deviation from $1-\eta$ in both low and high frequency ranges, and the overall trend appears distorted. In contrast, the TF24 array demonstrates a more linear relationship. The fact that the diffuseness does not reach exactly unity when $\eta=0$ (reverberation room only) and does not converge to zero when $\eta=1$ (anechoic room only) is attributed to both the non-ideal characteristics of the measured sound fields and the inevitable experimental errors, so the absolute values should be interpreted with appropriate caution.
Figure 14: The maximum absolute difference between the calculated diffuseness and $1-\eta$ for each index
As a reference, Fig. 14 shows the maximum absolute difference between the calculated diffuseness and $1-\eta$ for each index. The overall trends are consistent with those in Fig. 11: the intensity-energy (I/E) indices generally yield slightly larger errors than the eigenvalue-based indices, and the Afmt array exhibits larger deviations than the TF24 array.
VI Conclusion
This study presented a comparative investigation of diffuseness estimation methods based on intensity-energy relations and eigenvalue analysis, using several microphone arrays including the Afmt, Fibo64, and TF24 systems. Through a series of controlled simulations, the relationships between the diffuseness index and the beam-to-diffuse ratio were systematically examined under single-source, mixed, and interference conditions. The results demonstrated that the eigenvalue-based indices, particularly $\Psi_{\mathrm{COM}}$, exhibited higher robustness and a more linear response than the conventional intensity-energy (I/E) approach. The I/E method tended to overestimate diffuseness in interference fields where the active intensity approaches zero, leading to physically inconsistent interpretations.
Regarding the combination of microphone array configurations and evaluation approaches, the Fibo64 array would be the first choice if such an array is available. In cases where this is not realistic, arrays composed of microphones with known directivity patterns—such as the Tight frame configuration, which possesses theoretical validity—represent a reasonable alternative. Furthermore, the eigenvalue analysis based on the covariance matrix constructed only from the FOA velocity components ($u_{x}$, $u_{y}$, and $u_{z}$) was found to produce stable and reliable results overall.
Experimental verification using impulse responses measured in anechoic and reverberation rooms further confirmed the general trends predicted by the simulations. The diffuseness estimated from mixed impulse responses showed approximately linear dependence on the energy ratio between the direct and reverberant components, with smaller deviations observed for the TF24 array compared with the Afmt array. These results suggest that the velocity-based covariance analysis provides a practical and reliable framework for characterizing the spatial diffuseness of sound fields, even without strict whitening or mode-power normalization.
Future work will extend the present analysis beyond the weighted averaging of directional diffuseness, exploring how this parameter relates to perceptual impressions of spatial quality. Further investigations will aim to establish practical applications of diffuseness-based metrics in spatial reproduction and acoustic design, where such perceptually informed measures may provide useful guidance for sound field control and evaluation.
ACKNOWLEDGMENTS
The author would like to thank Mr. T. Ohno, Mr. R. Ohnuki, and Mr. K. Hirota, a graduate of his laboratory, for their contributions during the early stages of this research and measurements. The author also thanks Dr. H. Kashiwazaki for his valuable suggestions and insightful comments regarding the Tight frame structure. This research was supported by JSPS KAKENHI Grant No. JP24K03222.